# GraphCheck: Multi-Path Fact-Checking with Entity-Relationship Graphsthanks: Code available at: https://github.com/windowh1/graphcheck

### v2

@@ 3.4 DP-GraphCheck
Determine strategy $\sigma\leftarrow\mathcal{S}(C)$
if $\sigma=\textsc{Direct}$ then
@@
         if not $triplet\_supported$ then
@@
     if $path\_supported$ then
@@
To address this limitation, we introduce DP-GraphCheck, which improves both efficiency and accuracy. Given the claim and its top-$k$ retrieved documents (using the original claim itself as a query), a lightweight strategy selector determines whether the retrieved evidence is sufficient for assessing its veracity. If deemed sufficient, the claim is considered simple and verified via Direct. Otherwise, the claim undergoes the full GraphCheck pipeline, including claim decomposition and latent entity infilling.
This framework efficiently filters out simple claims, while maintaining thorough verification for more complex cases. The complete verification process of DP-GraphCheck is summarized in Algorithm 1.
4 Experiments
4.1 Experimental Setup
Datasets
We utilize two datasets:
  • HOVER Jiang et al. (2020) is a dataset for multi-hop fact-checking, verifying whether a claim is supported or not based on evidence dispersed across multiple Wikipedia articles (2-hop, 3-hop, 4-hop). Since the test set labels are not publicly available, we use the development set as our test set. We use the preprocessed October 2017 Wikipedia dump Yang et al. (2018) as the retrieval corpus.
  • EX-FEVER Aly et al. (2021) is another multi-hop fact-checking dataset, where evidence is scattered across multiple Wikipedia articles (2-hop, 3-hop). Unlike HOVER, which has only two labels, EX-FEVER introduces an additional ‘Not Enough Information (NEI)’ label. To align with the HOVER dataset, we exclude samples labeled as NEI. Following the dataset authors’ guidance, we use the same October 2017 Wikipedia dump as HOVER for the retrieval corpus.
Implementation Details
We employ FLAN-T5-XL Chung et al. (2022) for text infilling, fact triplet verification, strategy selection, and Direct, applying task-specific prompts (Appendix B). This setup aligns with our baseline, ProgramFC Pan et al. (2023b), ensuring a fair comparison.
For retrieval, we employ BM25 Robertson and Zaragoza (2009), maintaining consistency with ProgramFC as well. Our primary focus is the open-book setting, where the verification is conducted based on the top-$k$ retrieved documents ($k{=}10$). We also evaluate performance in the open-book + gold setting, where the claim’s gold document set is merged with the top retrieved documents to form a set of $k$ documents.
For entity-relationship graph construction, we employ gpt-4o (gpt-4o-2024-08-06). The few-shot examples used in the graph construction prompt are manually annotated using 10 instances randomly sampled from the HOVER training set (Appendix B.1).
Baselines
  • ProgramFC Pan et al. (2023b) is a fact-checking framework that decomposes complex claims into Python-like reasoning programs, outlining step-by-step actions such as question answering and subclaim verification. We generally follow the original experimental setup; however, since ProgramFC originally used Codex Chen et al. (2021), we re-run the experiments using gpt-4o to ensure comparability. We evaluate two cases, $N{=}1$ and $N{=}5$, where $N$ represents the number of OpenAI API calls, with each call expected to generate a distinct reasoning program.
  • FOLK Wang and Shu (2023) decomposes claims into First-Order Logic (FOL) clauses and question-answering sets required for fact-checking. Originally, FOLK used gpt-3.5 (text-davinci-003) for decomposing and SerpAPI for evidence retrieval and question answering. To align with our setup, we adapt FOLK to use gpt-4o and replace its question-answering module with FLAN-T5-XL, which generates answers based on Wikipedia articles retrieved via BM25.
  • FLAN-T5-XL Chung et al. (2022), a model with 3 billion parameters, has shown strong generalization and reasoning capabilities across various tasks. As it serves as the Direct component in DP-GraphCheck, we also report its standalone performance.
Additionally, we apply the strategy selector to ProgramFC and FOLK, to evaluate its generalizability.
Methods Open-book Open-book + Gold
HOVER EX-FEVER HOVER EX-FEVER
2-hop 3-hop 4-hop 2-hop 3-hop 2-hop 3-hop 4-hop 2-hop 3-hop
Direct (DP) 72.56 61.70 59.57 81.03 73.02 76.03 67.18 61.26 87.77 81.82
ProgramFC ($N{=}1$) 70.04 61.33 59.00 77.55 71.50 71.52 64.74 63.99 83.82 78.52
DP-ProgramFC ($N{=}1$) 70.79 62.75 60.61 79.46 74.62 71.19 66.04 65.31 83.42 78.93
ProgramFC ($N{=}5$) 70.29 61.82 60.19 78.31 72.16 70.73 65.50 63.79 84.09 79.69
DP-ProgramFC ($N{=}5$) 70.71 63.39 61.56 79.78 75.30 70.52 66.23 64.86 83.69 79.80
FOLK 65.13 59.63 56.10 72.32 63.85 69.96 66.23 65.65 80.13 75.98
DP-FOLK 70.96 63.53 58.36 80.64 73.68 72.48 69.39 68.59 84.99 80.66
GraphCheck ($\bar{P}{=}1$) 73.05 64.87 59.19 75.71 65.02 78.18 70.68 67.70 83.06 75.86
DP-GraphCheck ($\bar{P}{=}1$) 76.29 67.36 62.35 81.12 74.56 77.25 73.00 71.95 85.78 82.87
GraphCheck ($\bar{P}{=}5$) 74.12 67.71 64.79 76.56 69.94 78.59 73.78 72.55 83.64 80.16
DP-GraphCheck ($\bar{P}{=}5$) 76.29 68.70 66.64 81.12 76.02 76.96 73.34 73.63 85.69 82.73
Table 1: Macro-F1 scores under open-book and open-book + gold settings. The best result in each column is highlighted in bold, and the second-best result is underlined. The ‘DP-’ notation indicates that the strategy selector is used to combine Direct with the corresponding method.
4.2 Main Result
Table 1 summarizes the Macro-F1 scores for both the open-book and open-book + gold settings. The ‘DP-’ notation indicates that the strategy selector is used to combine Direct with the corresponding method.
DP-GraphCheck
In nearly all configurations—including the HOVER and EX-FEVER dataset, under both open-book and open-book + gold settings—DP-GraphCheck consistently achieves either the best or second-best Macro-F1 score. This demonstrates the strength of our framework, which combines the best of both worlds: it maintains the efficiency and precision of direct prompting for simple claims, while leveraging the structured reasoning of GraphCheck for more complex ones. As a result, DP-GraphCheck achieves the strongest overall performance.
An exception occurs in the EX-FEVER open-book + gold 2-hop setting, where the Direct method achieves the highest score. This can be attributed to EX-FEVER’s extractive nature, where claims closely match phrases in the gold documents with minimal rephrasing or abstraction (see Appendix A for a specific example). In such cases, Direct, which preserves the claim’s original form, may yield strong results. However, this advantage may not generalize to real-world fact-checking scenarios, where claims often differ in wording from the supporting evidence. In these realistic settings, more structured reasoning becomes essential.
GraphCheck
In the HOVER dataset, under both the open-book and open-book + gold settings, GraphCheck with $\bar{P}{=}5$ outperforms all baselines, even without the strategy selector. This highlights the effectiveness of graph-based structured verification in settings that require deeper reasoning and decomposition.
In the EX-FEVER dataset, GraphCheck occasionally underperforms compared to baseline methods. Considering the extractive nature of EX-FEVER, these results may be attributed to GraphCheck’s enforced fine-grained decomposition, in which the claim is broken down into a set of entity-relation triplets, each corresponding to a subclaim. While such structured decomposition is beneficial for complex reasoning, it may be unnecessary for extractive-style claims that can be verified holistically. However, when combined with Direct (i.e., DP-GraphCheck), GraphCheck consistently recovers strong performance.
Multi-path Verification
We also observe that GraphCheck with $\bar{P}{=}5$ generally outperforms its $\bar{P}{=}1$ counterpart, and this improvement frequently carries over to the DP variants. The performance gap becomes more pronounced as the hop count increases. For example, in the HOVER open-book setting, GraphCheck ($\bar{P}{=}5$) surpasses GraphCheck ($\bar{P}{=}1$) by 1.07, 2.84, and 5.60 points in 2-hop, 3-hop, and 4-hop claims, respectively. These results indicate that multi-path exploration becomes increasingly beneficial for more complex claims.
While ProgramFC also attempts multi-path verification by generating $N$ distinct verification programs via $N$ independent LLM API calls, the performance gap between ProgramFC ($N{=}1$) and ProgramFC ($N{=}5$) remains relatively small. This suggests that the reasoning paths generated by each call may lack sufficient diversity. Moreover, this method incurs high computational cost due to repeated LLM usage, even for claims that may not require extensive reasoning.
In contrast, GraphCheck constructs a single entity-relation graph through one LLM call, from which diverse verification paths can be explored. This design naturally promotes reasoning diversity without additional LLM overhead.
Effectiveness of Strategy Selector
The strategy selector is effective across all baselines: both DP-ProgramFC and DP-FOLK consistently outperform their original versions. These results show that the strategy selector not only strengthens our own framework, but also serves as a general enhancement applicable to existing fact-checking methods.
4.3 Ablation Study
Breakdown of DP-GraphCheck
Table 2 presents a detailed breakdown of DP-GraphCheck’s performance on the HOVER dataset under the open-book setting. It reports results for the entire dataset as well as for two subsets of claims, each assigned to either Direct or GraphCheck by the strategy selector. For each group, the table shows the proportion of claims, the recall of retrieval results using the original claim as the query (i.e., the evidence provided to both the strategy selector and Direct), and the verification accuracy of both Direct and GraphCheck.
We observe that the strategy selector assigns an increasing proportion of claims to GraphCheck as hop count increases: 59.68% of 2-hop claims are handled by GraphCheck, compared to 88.35% for 4-hop claims. This trend indicates that the strategy selector is capable of discerning claim difficulty and assigning complex cases to a more systematic verification method.
Moreover, claims verified via Direct exhibit consistently higher retrieval recall than those verified via GraphCheck. This suggests that the strategy selector effectively assesses whether retrieved evidence is sufficient to support Direct verification, thereby avoiding unnecessary decomposition.
In terms of accuracy, within the group of claims verified via Direct, the accuracy of the Direct consistently exceeds that of GraphCheck across all hop levels, demonstrating the effectiveness of direct prompting for simpler claims. Conversely, for claims verified via GraphCheck, the GraphCheck method outperforms Direct, particularly on 3-hop and 4-hop claims, highlighting the importance of structured reasoning in complex scenarios.
Overall, These results indicate that DP-GraphCheck successfully combines the strengths of Direct and GraphCheck through the strategy selector, achieving not only greater efficiency but also improved overall performance.
Group Metric 2-hop 3-hop 4-hop
Total Recall@10 73.18 51.34 36.43
Accuracy (Direct) 72.56 62.02 59.58
Accuracy (GraphCheck) 74.60 68.12 66.79
Verified via Direct % of Claims 40.32% 23.22% 11.65%
Recall@10 84.47 64.32 55.79
Accuracy (Direct) 74.01 71.13 72.73
Accuracy (GraphCheck) 69.82 68.54 64.46
Verified via GraphCheck % of Claims 59.68% 76.78% 88.35%
Recall@10 37.71 29.76 21.66
Accuracy (Direct) 71.58 59.26 57.84
Accuracy (GraphCheck) 71.48 67.99 67.10
Table 2: Breakdown of DP-GraphCheck performance on the HOVER dataset under the open-book setting. Results are grouped by the strategy selector’s assignment: total samples, samples verified via Direct, and those verified via GraphCheck. Bold values indicate the accuracy of the verification method actually used for each group.
Impact of Graph Construction Model
A core component of GraphCheck is its graph construction step, which requires decomposing a claim into fact triplets and detecting latent entities simultaneously. To assess the generalizability of GraphCheck across different language models beyond our default choice (gpt-4o), we experiment with several LLMs using identical prompts.
As shown in Table 3, smaller models such as gpt-3.5-turbo and Qwen2.5-72B-Instruct yield relatively lower performance in GraphCheck compared to gpt-4o, indicating that model capacity can influence the quality of graph construction. However, Qwen2.5-72B-Instruct, despite its relatively modest size, demonstrates surprisingly competitive results. Moreover, the performance gap between Qwen2.5-72B-Instruct and gpt-4o narrows considerably in DP-GraphCheck. For 2-hop claims, Qwen2.5-72B-Instruct achieves an even higher score (77.00 vs. 76.29), suggesting that the strategy selector effectively compensates for imperfect graph construction by leveraging direct prompting when appropriate.
In the case of claude-3-7-sonnet, the model demonstrates strong performance in both GraphCheck and DP-GraphCheck method, outperforming gpt-4o on 2-hop and 3-hop claims.
These findings suggest that the strong performance of GraphCheck and DP-GraphCheck with gpt-4o can be reproduced with other language models as well. While graph construction quality may vary depending on the underlying model—particularly for smaller LLMs—the DP mechanism helps compensate for suboptimal graphs by leveraging Direct when it is more appropriate. This demonstrates that GraphCheck and DP-GraphCheck are model-agnostic frameworks that generalize well across a range of LLM backbones.
Graph Construction Model Method 2-hop 3-hop 4-hop
gpt-4o-2024-08-06 (Default) GraphCheck 74.12 67.71 64.79
DP-GraphCheck 76.29 68.70 66.64
gpt-3.5-turbo GraphCheck 70.06 59.93 59.35
DP-GraphCheck 74.42 63.68 62.00
Qwen2.5-72B-Instruct GraphCheck 73.04 62.33 62.74
DP-GraphCheck 77.00 66.49 65.55
claude-3-7-sonnet-20250219 GraphCheck 76.39 68.53 63.65
DP-GraphCheck 76.45 69.74 65.50
Table 3: Macro-F1 scores of GraphCheck and DP-GraphCheck on the HOVER dataset (open-book setting) using different graph construction models.
Various Document-level Settings
We compare three document-level settings for verifying claims based on the top-$k$ retrieved documents: (1) concat, where the claim is verified once over the concatenation of all documents; (2) each, where the claim is evaluated against each document individually and considered Supported if any yields a Supported judgment; and (3) concat+each, which combines both approaches and returns Supported if either the concatenated context or any individual document supports the claim.
In DP-GraphCheck, Direct adopts the concat setting, while GraphCheck employs the concat+each setting. As shown in Table 4, these choices align well with the nature of input claim each method receives.
For Direct, the input is the original multi-hop claim, which typically integrates information distributed across multiple documents. Verifying the claim on each document in isolation may fail to capture this broader context, likely degrading performance. As a result, evaluating claims over the concatenated evidence (concat) shows the strongest performance.
In contrast, GraphCheck verifies fine-grained subclaims, each represented as an entity-relation triplet. These subclaims often correspond to atomic facts localized within individual documents, making the each setting more appropriate. However, some triplets still require cross-document context, which makes concat+each the most effective setting for GraphCheck.
These results show the importance of aligning the document-level setting with the nature of the input claim—whether it is a original multi-hop claim or a decomposed triplet—to achieve optimal verification performance.
Method Document-level Setting 2-hop 3-hop 4-hop
Direct concat 72.56 61.70 59.57
each 65.55 59.17 57.70
concat+each 67.90 59.92 58.40
GraphCheck concat 69.39 62.64 60.07
each 74.79 66.19 62.36
concat+each 74.12 67.71 64.79
Table 4: Macro-F1 scores of Direct and GraphCheck under different document-level settings on the HOVER dataset (open-book setting). Bold values indicate the actual document-level setting used by each method.
5 Conclusion
We introduce GraphCheck, a novel framework for automated fact-checking based on entity-relationship graphs, specifically designed to handle multi-hop reasoning. By converting claims into structured entity-relationship graphs and exploring multiple identification paths, our approach enables comprehensive and systematic verification. Additionally, we proposed DP-GraphCheck, which leverages a lightweight strategy selector to adaptively choose between simple direct prompting and more systematic graph-based verification, thereby enhancing both efficiency and robustness.
Our experiments on the HOVER and EX-FEVER datasets demonstrate that DP-GraphCheck consistently achieves superior performance across multiple settings. Furthermore, the strategy selector generalizes well to other fact-checking methods, making it a practical and adaptable component for broader use. Ablation studies further validate the effectiveness of each component, including multi-path traversal and document-level verification strategies. Overall, our findings highlight GraphCheck and DP-GraphCheck as strong and extensible frameworks for multi-hop claim verification.
Limitations
Despite its advantages, our framework has some limitations. First, the construction of entity-relationship graphs can be error-prone, which may propagate to the verification stage. Common issues include misclassifying non-latent entities as latent and failing to detect actual latent entities. Although this work focuses on a few-shot prompting setting, we believe that incorporating training could improve the quality of graph construction.
Second, our framework currently operates solely over a textual knowledge base. Given that the verification process is grounded in a structured graph representation, future extensions could explore applying GraphCheck in settings where the evidence is sourced from structured knowledge bases such as knowledge graphs.
Appendix A An Example from the EX-FEVER Dataset
EX-FEVER is highly extractive compared to the HOVER dataset. An example from EX-FEVER is shown in Figure 2. As illustrated in the example, claims in EX-FEVER closely match phrases in the gold documents with minimal rephrasing or abstraction.
@@ Appendix A An Example from the EX-FEVER Dataset
Appendix B Prompts
B.1 Graph Construction
@@
B.2 Latent Entity Infilling
@@
B.3 Direct (Fact Triplet Verification)

### v3

@@ Appendix E DP-GraphCheck Algorithm
Determine strategy $\sigma\leftarrow\mathcal{S}(C,d_{\text{concat}})$
if $\sigma=\textsc{Direct}$ then
@@
   if not triplet_supportedtriplet\_supported then
@@
  if path_supportedpath\_supported then
@@
Appendix F An Example from the EX-FEVER Dataset
@@
EX-FEVER is highly extractive compared to the HOVER dataset. As illustrated in the example shown in Figure 2, claims in EX-FEVER closely match phrases in the gold documents with minimal rephrasing or abstraction.
Appendix G Dataset Statistics and Model Specifications
G.1 Dataset Statistics
HOVER Jiang et al. (2020) consists of 18,171 training, 4,000 development, and 4,000 test claims, requiring 2 to 4 hops of reasoning. Labels are Supported or Not Supported. Since test labels are not publicly available, we use the development set for evaluation.
EX-FEVER Ma et al. (2024) contains 43,107 training, 12,059 development, and 6,099 test claims, with 2 to 3 hop reasoning. Labels are Supports, Refutes, and Not Enough Info (NEI). Based on discussions with the dataset creators, we confirm that the NEI label does not always indicate a true absence of evidence in the retrieval corpus, but only in the annotated subset. In open-domain retrieval settings, NEI-labeled claims can often be verified with evidence, making NEI label less reliable for evaluation. We therefore exclude NEI samples from our experiments.
FEVEROUS Aly et al. (2021) consists of 71,291 training, 7,890 development, and 7,845 test claims. Labels are Supported, Refuted, and NEI. FEVEROUS involves evidence from both unstructured text and structured tables. Since we focus on textual fact-checking, we adopt the subset of Pan et al. (2023b), which selects only claims that require sentence-only evidence from the development set, yielding 2,962 claims for evaluation.
FEVER Thorne et al. (2018) is a large-scale single-hop fact-checking dataset with 145,449 training, 9,999 development, and 9,999 test claims. Labels are Supported, Refuted, and NEI. The dataset consists mostly of simple claims that require single-hop reasoning, such as “The Beatles was a rock band”.
G.2 Model Specifications
We experiment with a range of LLMs of different scales and access types.
For open-source models, we use flan-t5-xl (3B parameters) and Qwen2.5-72B-Instruct (72B parameters), both released on Hugging Face.
Among API-based proprietary models, we use gpt-4o-2024-08-06 and gpt-3.5-turbo from OpenAI, whose parameter counts and architectural details have not been publicly disclosed. We further include claude-3.7-sonnet-20250219 from Anthropic, also without public parameter specifications.
G.3 Licenses and Terms of Use
HOVER is released under the CC BY-SA 4.0 license, while EX-FEVER is provided for research purposes. FEVER is distributed under the CC BY-SA 3.0 and GPL-3.0 licenses, and FEVEROUS is released under the CC BY-SA 4.0 license. flan-t5-xl and Qwen2.5-72B-Instruct are available on Hugging Face under the Apache 2.0 and Qwen license, respectively. API-based proprietary models (gpt-4o-2024-08-06, gpt-3.5-turbo, and claude-3.7-sonnet-20250219) are accessed in accordance with their respective terms of use. All artifacts are used strictly for research purposes and in a manner consistent with their intended use and license conditions.
Appendix H Prompts
H.1 Graph Construction
@@
H.2 Latent Entity Infilling
@@
H.3 Direct (Fact Triplet Verification)
@@ H.4 Strategy Selection
Appendix I Examples of Graph Construction Error
(1) Hallucinated Latent Entities # Claim:
 Age of Chance and Vector were both a rock band
# Latent Entities:
 (ENT1) [SEP] is [SEP] a rock band
 (ENT2) [SEP] is [SEP] a rock band
# Triples:
 Age of Chance [SEP] is [SEP] (ENT1)
 Vector [SEP] is [SEP] (ENT2) Issue: Non-latent entities (“Age of Chance”, “Vector”) are mistakenly marked as latent, leading to unnecessary and potentially erroneous infilling.
(2) Decomposition Error # Claim:
 The place where MLB team owner Doc Watson (baseball) was born is named after the only Catholic signatory to the Declaration of Independence, Charles Carroll of Carrollton.
# Latent Entities:
 (ENT1) [SEP] is [SEP] a place
# Triples:
 Doc Watson (baseball) [SEP] was born in [SEP] (ENT1)
 (ENT1) [SEP] is named after [SEP] Charles Carroll of Carrollton
 Charles Carroll of Carrollton [SEP] is [SEP] the only Catholic signatory to the Declaration of Independence Issue: Important context (“MLB team owner”) is dropped during decomposition.
Table 16: Examples of graph construction error in GraphCheck.